# Understanding Multimodal LLMs Under Distribution Shifts: An Information-Theoretic Approach

### v1

@@ D.2 Proof for EMID upper bound
$\begin{split}H(P_{Y|\mathbf{X}})-H(Q_{Y|\mathbf{X}})&\leq H(Q_{Y})\sqrt{D_{\rm JS}(P_{X_{v}}||Q_{X_{v}})+D_{\rm JS}(P_{X_{t}}||Q_{X_{t}})+\bar{D}_{\rm JS}(P_{X_{t}|X_{v}}||Q_{X_{t}|X_{v}})+\bar{D}_{\rm JS}(P_{X_{v}|X_{t}}||Q_{X_{v}|X_{t}})}\\ &+4\mathbb{E}_{P_{\mathbf{X}}}[D^{\frac{1}{4}}_{\rm JS}(P_{Y|\mathbf{X}=\mathbf{x}}||Q_{Y|\mathbf{X}=\mathbf{x}})]\\ &\leq\hat{H}(Q_{Y|\mathbf{x}})\big{(}D^{\frac{1}{2}}_{\rm JS}(P_{X_{v}}||Q_{X_{v}})+D^{\frac{1}{2}}_{\rm JS}(P_{X_{t}}||Q_{X_{t}})\big{)}\\ &+\hat{H}(Q_{Y|\mathbf{x}})\big{(}\bar{D}^{\frac{1}{2}}_{\rm JS}(P_{X_{t}|X_{v}}||Q_{X_{t}|X_{v}})+\bar{D}^{\frac{1}{2}}_{\rm JS}(P_{X_{v}|X_{t}}||Q_{X_{v}|X_{t}})\big{)}\\ &+4\mathbb{E}_{P_{\mathbf{X}}}[D^{\frac{1}{4}}_{\rm JS}(P_{Y|\mathbf{X}=\mathbf{x}}||Q_{Y|\mathbf{X}=\mathbf{x}})].\end{split}$
@@
Finally, we complete the proof by adding the ineq. (D.2) and ineq. (19) to induces the upper bound of $\text{EMID}(P_{\mathbf{X}Y},Q_{\mathbf{X}Y};\theta)$,
$\begin{split}&\text{EMID}(P_{\mathbf{X}Y},Q_{\mathbf{X}Y};\theta)\\ &\leq(H(P_{Y|\mathbf{X}})-H(Q_{Y|\mathbf{X}}))+\big{(}\mathbb{E}_{\mathbf{x}\sim Q_{\mathbf{X}}}[H(P_{\theta}(\cdot|\mathbf{x}))]-\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[H(P_{\theta}(\cdot|\mathbf{x}))]\big{)}+8\Delta^{\frac{1}{4}}\\ &\leq\hat{H}\big{(}D^{\frac{1}{2}}_{\rm JS}(P_{X_{v}}||Q_{X_{v}})+D^{\frac{1}{2}}_{\rm JS}(P_{X_{t}}||Q_{X_{t}})\big{)}\\ &+\hat{H}\big{(}\bar{D}^{\frac{1}{2}}_{\rm JS}(P_{X_{t}|X_{v}}||Q_{X_{t}|X_{v}})+\bar{D}^{\frac{1}{2}}_{\rm JS}(P_{X_{v}|X_{t}}||Q_{X_{v}|X_{t}})\big{)}\\ &+4\mathbb{E}_{P_{\mathbf{X}}}[D^{\frac{1}{4}}_{\rm JS}(P_{Y|\mathbf{X}=\mathbf{x}}||Q_{Y|\mathbf{X}=\mathbf{x}})]+8\Delta^{\frac{1}{4}}\end{split}$
@@
where $P_{Y_{\theta}}=\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}P_{{\theta}}(\cdot|\mathbf{x})$ and $Q_{Y_{\theta}}=\mathbb{E}_{\mathbf{x}\sim Q_{\mathbf{X}}}P_{{\theta}}(\cdot|\mathbf{x})$, then $\text{EMID}(P_{\mathbf{X}Y},Q_{\mathbf{X}Y};\theta)$ is upper bounded by
@@
Given an MLLM $P_{\theta}$ and distributions $P_{\mathbf{X}Y}$, $Q_{\mathbf{X}Y}$ which have consistent conditional distributions over variables $X_{v}|X_{t}$, $X_{t}|X_{v}$, and $Y|\mathbf{X}$, $\text{EMID}(P_{\mathbf{X}Y},Q_{\mathbf{X}Y};\theta)$ is upper bounded by

### v2

@@ D.2 Proof for EMID upper bound
$\begin{split}&H(P_{Y|\mathbf{X}})-H(Q_{Y|\mathbf{X}})\\ &\leq H(Q_{Y})\sqrt{D_{\rm JS}(P_{X_{v}}||Q_{X_{v}})+D_{\rm JS}(P_{X_{t}}||Q_{X_{t}})+\bar{D}_{\rm JS}(P_{X_{t}|X_{v}}||Q_{X_{t}|X_{v}})+\bar{D}_{\rm JS}(P_{X_{v}|X_{t}}||Q_{X_{v}|X_{t}})}\\ &+4\mathbb{E}_{P_{\mathbf{X}}}[D^{\frac{1}{4}}_{\rm JS}(P_{Y|\mathbf{X}=\mathbf{x}}||Q_{Y|\mathbf{X}=\mathbf{x}})]\\ &\leq\hat{H}(Q_{Y|\mathbf{x}})\big{(}D^{\frac{1}{2}}_{\rm JS}(P_{X_{v}}||Q_{X_{v}})+D^{\frac{1}{2}}_{\rm JS}(P_{X_{t}}||Q_{X_{t}})\big{)}\\ &+\hat{H}(Q_{Y|\mathbf{x}})\big{(}\bar{D}^{\frac{1}{2}}_{\rm JS}(P_{X_{t}|X_{v}}||Q_{X_{t}|X_{v}})+\bar{D}^{\frac{1}{2}}_{\rm JS}(P_{X_{v}|X_{t}}||Q_{X_{v}|X_{t}})\big{)}\\ &+4\mathbb{E}_{P_{\mathbf{X}}}[D^{\frac{1}{4}}_{\rm JS}(P_{Y|\mathbf{X}=\mathbf{x}}||Q_{Y|\mathbf{X}=\mathbf{x}})].\end{split}$
@@
Finally, we complete the proof by adding the ineq. (D.2) and ineq. (21) to induces the upper bound of $\text{EMID}(P_{\mathbf{X}Y},Q_{\mathbf{X}Y};P_{\theta})$,
$\begin{split}&\text{EMID}(P_{\mathbf{X}Y},Q_{\mathbf{X}Y};P_{\theta})\\ &\leq(H(P_{Y|\mathbf{X}})-H(Q_{Y|\mathbf{X}}))+\big{(}\mathbb{E}_{\mathbf{x}\sim Q_{\mathbf{X}}}[H(P_{\theta}(\cdot|\mathbf{x}))]-\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[H(P_{\theta}(\cdot|\mathbf{x}))]\big{)}+8\Delta^{\frac{1}{4}}\\ &\leq\hat{H}\big{(}D^{\frac{1}{2}}_{\rm JS}(P_{X_{v}}||Q_{X_{v}})+D^{\frac{1}{2}}_{\rm JS}(P_{X_{t}}||Q_{X_{t}})\big{)}\\ &+\hat{H}\big{(}\bar{D}^{\frac{1}{2}}_{\rm JS}(P_{X_{t}|X_{v}}||Q_{X_{t}|X_{v}})+\bar{D}^{\frac{1}{2}}_{\rm JS}(P_{X_{v}|X_{t}}||Q_{X_{v}|X_{t}})\big{)}\\ &+4\mathbb{E}_{P_{\mathbf{X}}}[D^{\frac{1}{4}}_{\rm JS}(P_{Y|\mathbf{X}=\mathbf{x}}||Q_{Y|\mathbf{X}=\mathbf{x}})]+8\Delta^{\frac{1}{4}}\end{split}$
@@
where $P_{Y_{\theta}}=\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}P_{{\theta}}(\cdot|\mathbf{x})$ and $Q_{Y_{\theta}}=\mathbb{E}_{\mathbf{x}\sim Q_{\mathbf{X}}}P_{{\theta}}(\cdot|\mathbf{x})$, then $\text{EMID}(P_{\mathbf{X}Y},Q_{\mathbf{X}Y};P_{\theta})$ is upper bounded by
@@
Given an MLLM $P_{\theta}$ and distributions $P_{\mathbf{X}Y}$, $Q_{\mathbf{X}Y}$ which have consistent conditional distributions over variables $X_{v}|X_{t}$, $X_{t}|X_{v}$, and $Y|\mathbf{X}$, $\text{EMID}(P_{\mathbf{X}Y},Q_{\mathbf{X}Y};P_{\theta})$ is upper bounded by
@@
D.3 Discussion
We made the $\epsilon$-representation capacity assumption to derive Lemma 4.3 and Theorem 4.4 that claim the theoretical justification for EMI by showing its connection to a classic preference model. The assumption captures a minimum achievable discrepancy between the truth distribution $P_{Y|\mathbf{X}}$ and the model’s distribution $P_{\theta}(\cdot|\mathbf{x})$. As the models become more expressive–e.g., by increasing model size (Kaplan et al., 2020) and leveraging advanced positional encoding (Luo et al., 2022)–MLLM approaches the universal approximator of sequence-to-sequence mapping (Luo et al., 2022; Furuya et al., 2024), as a result, the minimum expected discrepancy tends to decrease, leading to a smaller $\epsilon$.
Meanwhile, EMI and EMID trigger flexible potential use cases given their generality. For example, they can be used as metrics to evaluate a pure LLM as well as a multimodal LLM, depending on the types of targeted problems. Moreover, although we confined our analysis on $I(P_{\mathbf{X}Y})$, the chain rule of the mutual information (Cover, 1999) allows us to conduct partial modality EMI analysis through $I(P_{\mathbf{X}Y})=0.5I(P_{X_{v}Y})+0.5I(P_{X_{t}Y})+0.5I(P_{X_{t}Y|X_{v}})+0.5I(P_{X_{v}Y|X_{t}})$, where we can decompose the aggregated EMI into the modality-specific EMI terms and modality interaction (conditional) EMI terms.